Solve the following equations. 
6 = z + 2
Solve the following equations. 
6 = z + 2

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

6=z+2
z+2=6
z = 6-2
z=4

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
6 = z + 2
Z = 6 - 2
Z = 4
 
Hope it helps !
Thank you ✌️